### v1

# CommVQA: Situating Visual Question Answering in Communicative Contexts

Abstract
Current visual question answering (VQA) models tend to be trained and evaluated on image-question pairs in isolation. However, the questions people ask are dependent on their informational needs and prior knowledge about the image content. To evaluate how situating images within naturalistic contexts shapes visual questions, we introduce CommVQA, a VQA dataset consisting of images, image descriptions, real-world communicative scenarios where the image might appear (e.g., a travel website), and follow-up questions and answers conditioned on the scenario. We show that CommVQA poses a challenge for current models. Providing contextual information to VQA models improves performance broadly, highlighting the relevance of situating systems within a communicative scenario.
1 Introduction
Visual question answering (VQA), the task of providing an answer to a question given an image, measures a model’s ability to synthesize visual and textual modalities, and has many promising real-world applications. For example, when images online can’t be seen and accessed, it severely affects people’s abilities to educate themselves, socially engage, and stay informed Morris et al. (2016); MacLeod et al. (2017); Voykinska et al. (2016); Gleason et al. (2019), and VQA models are an opportunity for providing interactive accessibility to such visual content at scale Gurari et al. (2018); Baker et al. (2021). While most prior VQA datasets focus on investigating image-text alignment as a decontextualized task Antol et al. (2015); Goyal et al. (2017); Hudson and Manning (2019); Marino et al. (2019), we aim to reframe it as a human-centric communicative problem, moving it closer to a real-world interactive setting.
According to a pragmatic Bayesian view of communicative actions, people tend to ask questions that maximize the contextually relevant information gain based on their existing prior beliefs about the world Frank and Goodman (2012), following fundamental pragmatic principles Grice (1975). Based on these linguistic insights, we argue that prior VQA datasets largely do not consider two central communicative drives that limit their utility. The information people aim to obtain (and consequently the types of questions they ask) varies with (1) the person’s information needs based on their goals when encountering the image, and (2) the person’s prior knowledge of the image content. Thus we introduce CommVQA, a benchmark that treats VQA as an inherently communicative task.  [^1] All data and code will be made available at: https://github.com/nnaik39/commvqa.
Figure 1: Experimental design overview. First, images were crawled from Wikipedia and matched with plausible context scenarios. Descriptions for each image were initially elicited from GPT-4V and then edited by humans to ensure truthful and useful image descriptions. Then questions and answers were elicited from people, conditioned on the context and the image description, resulting in a minimum of three answers to each of the 1,215 unique visual questions. Instructions are simplified for illustration purposes (see Appendix A.1 for full prompts and task setups).
To investigate people’s information needs, CommVQA consists not only of images, questions, and answers, but also of image descriptions and plausible communicative scenarios for each image. Prior image accessibility research with blind and low-vision (BLV) participants shows that the information people want from an image is dependent upon the scenario in which the image appears Stangl et al. (2021, 2020); Kreiss et al. (2022a); Muehlbradt and Kane (2022). For example, if a person encounters an image when they are shopping online, they are likely to ask questions about the brands within the image, while if they’re browsing the news, the perceived purpose of the image shifts, and they are likely to ask questions about the event occurring within the image Stangl et al. (2021, 2020). We therefore define a scenario as the type of website (e.g., a shopping website) combined with a goal for viewing it (e.g., to buy a gift) and expect it to shape people’s information needs in a VQA task.
In addition, the relevant questions a person might ask are predicted to be guided by their prior knowledge about the image. For most images we encounter online, we can rely on rich cues that allow sophisticated inferences about what an image contains. An image on a shopping website, for instance, would likely be accompanied by an article label, such as “Colorful Summer Skirt”, or it could have an informative alt text description. In the standard VQA task, annotators are asked to write questions in isolation to “fool a smart robot”, an adversarial task where the goal of the questioner is to trick an AI model Goyal et al. (2017). However, in the naturalistic setting, these visual questions are better conceptualized as follow-up questions, since they are conditioned on already available information. In CommVQA, we situate the task by providing people with quality-controlled image descriptions instead of the image itself when collecting visual questions. Together with the contextual grounding of the images, this pipeline, presented in Figure 1, generates a challenging dataset with context-sensitive, highly diverse questions, as well as longer answers compared to prior VQA datasets.
We benchmark four state-of-the-art VQA models on CommVQA, and show that the situated nature of CommVQA poses a significant challenge for current state-of-the-art vision-language models (VLMs) and allows for analytical insights into model generation behavior due to a highly controlled data generation pipeline. Crucially, we find that models benefit from accessing the contextual information, further suggesting that these are important resources that shape VQA in communicative settings.
In summary, our main contributions are: (a) we introduce CommVQA, a benchmark consisting of images, contexts, descriptions, questions, and answers, where models must infer the details most relevant to the questioner’s goals, (b) we benchmark this dataset on current VQA approaches and explore whether models can be instructed to integrate this contextual condition, and (c) we analyze model performance on this task.
2 Related Work
With CommVQA, we aim to situate the abstract VQA task within communicative settings. This builds on prior VQA dataset work (Section 2.1), and communicative insights from linguistics and human-computer interaction (Section 2.2)
2.1 VQA Datasets
Image-Question-Answer Input
Most VQA datasets focus on image-question-answer triplets, which are constructed in isolation from the real-world contexts in which the images might appear Antol et al. (2015); Goyal et al. (2017); Hudson and Manning (2019); Marino et al. (2019). The VizWiz dataset Gurari et al. (2018) stands out among VQA datasets as it focuses on real-world accessibility. The images in VizWiz are elicited from BLV users uploaded a picture to a phone app with a question about their real-world environment. The goals behind the questions and images is specific to BLV people exploring their physical environment. CommVQA is complementary to VizWiz, since it investigates how contextually situating images online affects a model’s VQA task performance.
Image-Question-Answer-Context Input
Some VQA datasets have explored integrating supplementary information to the image. Visual Dialog Das et al. (2017) aims to create visual chatbots, which can answer a question based on an image and the prior dialogic context. Similarly to CommVQA, the questions in Visual Dialog can be similarly conceptualized as follow-up questions based on an image description. However, the images, questions, and answers are still decontextualized from a specific information goal beyond the questioner wanting to understand the image. CommVQA instead extends the focus on varying the questioner’s broader goals in understanding the image.
The ScienceQA Lu et al. (2022) and VQAOnline Chen et al. (2023) datasets were constructed by crawling online resources. ScienceQA Lu et al. (2022) consists of multimodal science understanding questions, where each multiple choice question and answer is associated with a lecture and explanation, functioning as the context. VQAOnline Chen et al. (2023) consists of images, questions, and context sourced from StackExchange. The question is the title and the context is the body of the StackExchange post, which constitutes supplementary information that’s useful for answering the question. The notion of context in these datasets is limited to texts that provide supplementary information, whereas we specifically focus on the effect of changing scenarios on the VQA task. Additionally, due to the controlled generation pipeline, CommVQA allows for insights on how models perform under varying contextual conditions.
2.2 VQA as Communication
Within pragmatics, there is a general consensus that questions are grounded in contexts and sensitive to the goals of the interlocutors (Groenendijk and Stokhof, 1984; Ginzburg, 1996; Roberts, 1996; van Rooy, 2003). In line with this prediction, prior human-subject studies with BLV participants show that context influences the information that people want about an image Stangl et al. (2021, 2020); Muehlbradt and Kane (2022); Kreiss et al. (2022a). Stangl et al. (2021) find that on social media, they wanted to know more about the people and the activity of the person who posted the image, while if people were on a shopping website to purchase a gift for a friend, they expressed a desire to learn more about the objects within the image. Inspired by Stangl et al. (2021), we utilize a similar type of scenario. With CommVQA, we constitute a large-scale dataset where context-sensitivity emerges with sighted participants in question and answer behavior and can be studied at scale.
3 The CommVQA Dataset
The CommVQA dataset was constructed in five main steps, as visualized in Figure 1. First, we sourced images from Wikipedia, and elicited both plausible scenarios and descriptions for the images from GPT-4V OpenAI (2023a). To ensure description quality, we conducted a human-subject study to edit the descriptions. We then elicited questions and answers for the dataset from US-based crowdworkers on Prolific. All human-subject studies were conducted with IRB approval.
3.1 Image-Scenario Matching
We extracted 150 images from Wikipedia pages of topics related to at least one of the scenario conditions. Since each image appears in two scenarios and each image-scenario pair is annotated with at minimum four questions and each question with three answers, this results in 3,642 unique question-answer pairs. We prioritized high coverage of individual datapoints over breadth since the focus of this work is on uncovering the diverse contextual effects on the VQA task.
We chose six potential scenarios: Shopping, Travel, Science Magazines, News, Health, and Social Media. These scenarios were informed by prior work that showed people’s information needs varied across these scenarios Stangl et al. (2021), and allowed for an overlap between the images that appear in each scenario–for instance, images in travel blogs (e.g., a picture of a waterfall) could plausibly appear in an online science magazine.
To assign plausible scenarios to the image, we instructed GPT-4V OpenAI (2023a) to rank the six scenarios in order of descending plausibility, using the following prompt, which was inspired by the framing in prior work Stangl et al. (2021): “Imagine you are a person browsing the Internet. Please rank the following scenarios in which you might encounter this image: 1) You are browsing a shopping website, with the goal of purchasing an item or experience…” (The full prompt to GPT4-V is included in Appendix A.1.) Using a human-subject pilot study, we established that GPT-4V made reasonable choices for contexts, which enabled us to easily scale. From the top-three scenarios for each image, two scenarios were selected in order to balance out the co-occurrence of contexts.
3.2 Image Description Elicitation
We elicited descriptions for all images in our dataset. They form the basis for the follow-up questions participants asked, simulating the effect of someone encountering an alt text associated with the image online. Importantly, descriptions were generated out-of-context so we could analyze context effects on the questions and answers without the description being a confounding factor.
Automatic Description Draft Elicitation
We elicited initial description drafts by prompting GPT4-V with the phrase: Describe this image in less than 150 characters to someone who cannot see it. The length constraint follows a commonly-issued guide on best-practices for accessibility alt text writing OSU (2024). We chose Wikipedia for sourcing images due to the copyright permissions and image variety.
Description Editing
To ensure description quality, we conducted a human-subject study where participants edited the descriptions generated by GPT-4V OpenAI (2023a). These edits were intended to help balance for potential inaccuracies, possible misalignments with human description preferences, and the current design choices of GPT-4V. For instance, as of October 2023, the model refrains from explicitly identifying the people within an image OpenAI (2023b), even though proper names can be an important detail for a useful image description Bennett et al. (2021); MacLeod et al. (2017).
Each participant was shown six randomized trials with an image and description. They were instructed to edit the image description to correct any errors and make the it more useful to someone who cannot see it. Importantly, participants were not shown the context that each image was placed in to control for question variation across contexts. Participants could also choose to skip if no edits were needed. We recruited 44 participants and compensated them at the rate of \$13.50/hr.
We collected three description edits for each image description draft, and selected a random edited description for the final description, as long as the Jaccard distance between the original GPT-4V description and the edited description was less than $0.2$, to ensure that we weren’t selecting outliers.
3.3 Question Elicitation
To elicit visual questions, we recruited 139 participants, who were paid \$13.50/hr with an average completion time of seven minutes. In each trial, participants were given two pieces of information: an image description (e.g., “A group of people of various ages walking along a grassy path, with trees on one side.”), and a scenario for the image (e.g., Imagine you are browsing a Health website, with the goal of learning how to live a healthier lifestyle). Crucially, participants didn’t see the image to avoid priming for specific questions and simulate the visual inaccessibility of the image in the real-world scenario. Participants rated how likely the image would appear within the provided scenario, and were prompted to ask two questions they would like answered by someone who can see the image.
Through this setup, we elicited 1,569 questions in total, with at least four questions for each image-scenario pair.
3.4 Answer Elicitation
We elicited answers from 480 participants, who were paid \$13.50/hr, with an average completion time of seven and a half minutes. Participants were shown the image, question, context, and description. They were told that another user asked the question based on this image description, and asked to write an answer that would help the other person visualize the image. In the final dataset, each question was answered by at least three separate participants. We also included a checkbox for participants to indicate if a question was unanswerable. If a question was voted unanswerable by two or more annotators, we excluded the question from the dataset, resulting in 355 exclusions (22.6%).
4 Dataset Analysis
The final CommVQA dataset consists of 3,642 unique question-answer pairs, spanning across 300 unique image-scenario pairs. To better understand the challenges posed by this dataset, we now provide an analysis of CommVQA. Examples from the dataset, sampled to cover all scenarios, are given in Appendix Figure 6.
4.1 Analysis of Descriptions
In CommVQA, the descriptions are the basis for the VQA task and were collected in a two-stage process: first an automatic description generation by GPT-4V and then a human editing phase. The automatically generated descriptions had an average length of $99.36$ characters and in $50\%$ of trials, people didn’t make any edits to the descriptions. Additions had a minimal effect on length–when participants chose to make edits, they added extra information, resulting in an average length of $106.04$ characters.
4.2 Analysis of Questions
Figure 2: Heatmap displaying the accuracies that finetuned BERT achieved for classifying questions into each pair of scenarios. BERT achieved an accuracy of $91.5\%$ for distinguishing science magazines and shopping, yet BERT only achieved a performance of $86\%$ to distinguish questions between science magazines and news.
The main goal of CommVQA is to situate the VQA task in a communicative context, assuming that the context shapes what becomes relevant and therefore has implications for downstream model performance. In this section, we investigate the context-sensitivity of the questions in CommVQA.
If questions are context-sensitive, then a trained classifier should be able to predict the correct context from the question and achieve a performance reliably higher than random. To investigate whether this is true for CommVQA, we fine-tuned BERT Devlin et al. (2019) on the task of predicting whether each question appears within a certain context or not. We expect this task to be difficult even for a human, given that questions such as “What time of day is it?” might appear in any context. We split all the questions collected into 20 cross-validation folds, with an 80-10-10 train/test/val split, and fine-tuned for 200 epochs with LoRA Hu et al. (2021). If questions written in different contexts do not differ, we should achieve only a random performance at $16\%$. However, fine-tuned BERT achieved an accuracy of $57\%$, a significant improvement over random choice, showing that questions elicited within different contexts are distinguishable even from a language-only perspective.
While we find overall evidence that questions vary between all contexts, a classifier analysis further allows us to investigate which individual contexts have the least and most overlap in the questions asked. Figure 2 shows BERT’s performance when fine-tuned on distinguishing pairs of scenarios (e.g., shopping from science magazines), and indicates which scenarios are more easily distinguishable. Intuitively, certain contexts are more related than others–for instance, shopping and social media content is more likely to contain images of famous models, while shopping and science magazines are less likely to have overlap. Our analysis confirms this intuition and highlights where the translational accuracy of models might fail downstream.
We further inspect the question interrogatives for a potential source of between-context variation. The results of a Welch’s t-test demonstrated that, for example, compared to all other scenarios, “Who” questions are significantly more likely to appear in the social media scenario ($t(310)=2.86$, $p<0.001$), and “What” questions are significantly more likely to appear in the shopping scenario ($t(388)=2.86$, $p<0.005$) compared to any other scenario. These results indicate that different question types are more likely to be asked in certain scenarios, which carries implications that model evaluation should be contextual. A model that has poor performance on “Who” questions might seem competent in non-social media scenarios, but fail to generalize to social media due to the distinct nature of the user’s information needs.
Taken together, we find converging evidence that the questions in CommVQA fundamentally vary based on the scenario the images were presented in, highlighting the diverse requirements for building robust communicatively situated VQA models.
4.3 Analysis of Answers
We now turn to a general analysis of the answers in CommVQA and investigate the extent to which they are contextually situated.
Firstly, with an average length of $9.8$ words, the answers in CommVQA are surprisingly long compared to prior datasets that used a similar online answer elicitation setup (e.g., 1.1 words for VQA-v2 Goyal et al. (2017) and 2.1 words for Visual Dialog Das et al. (2017)). Contrasting prior work, the instructions of CommVQA are framed as a communicative task, where answerers were asked to help someone else who cannot see the image. It’s therefore plausible that the longer responses are partially due to the participant’s wish to faithfully communicate with the questioner Grice (1975). In support of this hypothesis, we find that participants qualified many assumptions of theirs that were not directly supported by visual evidence with hedges and expressions of uncertainty, therefore lengthening the responses. In fact, the most common trigram within the dataset was “appear(s) to be” (93 occurrences), followed by “it looks like” (45 occurrences).
While we find that the scenario significantly shapes the content of the questions in CommVQA (see Section 4.2), it’s unclear to what extent we should expect the answers to be shaped by the scenario, as well. Since the questions are inherently affected by the context, it is difficult to establish a similar variation in the answers that isn’t already induced by the questions themselves. For example, we find that the rates at which named entities are mentioned largely varies between scenarios (see Appendix Figure 5), which is in line with the varied frequencies of “Who” questions. However, when we investigate the type of entities named in responses, we find additional effects of the scenario. For example, organizations are more frequently referenced in the travel scenario but individual people are more referenced in the science journal scenario (see Figure 3). This is borne out as a significant interaction between the proportion of mentioned organizations vs. people and the travel vs. science journal scenarios ($\beta=-2.22$, $SE=0.67$, $p<0.001$). This suggests that the external knowledge VQA models need to reference varies across scenarios and provides evidence that contextual considerations matter on the answer level.
Figure 3: Named entity recognition analysis showing significant variation within people’s answers between Travel and Science Journal scenarios. Organizations (e.g., companies, and institutions) were significantly more likely to be mentioned in the Travel over the Science scenario, and vice versa for people (e.g.., political figures, and celebrities). Asterisks indicate significant interaction. Error bars are 95% confidence intervals.
5 Model Benchmarking
Model BLEU-1 BLEU-4 METEOR ROUGE CIDEr CLIPScore
IDEFICS 0.326 0.106 0.189 0.401 0.825 0.667
IDEFICS (Contextual) 0.356 0.122 0.198 0.413 0.904 0.666
mPLUG-Owl 0.191 0.045 0.166 0.311 0.414 0.720
mPLUG-Owl (Contextual) 0.178 0.043 0.169 0.284 0.335 0.756
LLaVA 0.192 0.050 0.171 0.311 0.431 0.729
LLaVA (Contextual) 0.149 0.037 0.161 0.247 0.200 0.784
BLIP-2 0.289 0.067 0.100 0.286 0.435 0.604
BLIP-2 (Contextual) 0.059 0.010 0.078 0.224 0.343 0.662
Table 1: Results for the baseline and contextual condition across all models. The highest-scoring models in all columns are the ones which integrate the contextual condition. The CLIPScore metric improves for all models, except for IDEFICS, where it stays the same. This indicates the model’s generated answers contain more visual details with the contextual condition. Results were conducted over three confidence intervals.
In this section, we investigate the performance of four state-of-the-art VLMs on CommVQA  and to what extent providing context to the models improves their performance.
We selected four models to benchmark on this dataset: LLaVA Liu et al. (2023), BLIP-2 Li et al. (2023), mPLUG-OWL Ye et al. (2023), and IDEFICS Laurençon et al. (2023). To maximally enable the reproducibility of this work and support the public development of models, we focus on open-source models. LLaVA is trained on a multimodal instruction-following dataset, and exhibits strong performance across several VQA datasets Goyal et al. (2017); Hudson and Manning (2019). mPLUG-Owl Ye et al. (2023) is also instruction-tuned, and displays competitive performance on various VQA benchmarks, including VQA-v2 Goyal et al. (2017). IDEFICS Laurençon et al. (2023) is an open-access reproduction of Flamingo Alayrac et al. (2022), and was trained on a naturalistic web-scale dataset of interleaved image-text documents, including 141 million web pages. We selected LLaVA, mPLUG-Owl, and IDEFICS since their instruction-tuned nature allows us to straightforwardly integrate context. We also included a non-instruction-tuned model, BLIP-2 Li et al. (2023), a general-purpose vision-language model that is additionally fine-tuned on the VQA task.
Evaluation Metrics
We evaluated these models on the NLG metrics BLEU Papineni et al. (2002), CIDEr Vedantam et al. (2015), METEOR Banerjee and Lavie (2005), and ROUGE Lin (2004). However, there are still many aspects of performance these metrics cannot capture. For instance, studies show that these metrics do not adequately capture object hallucinations Rohrbach et al. (2018). To measure how closely an answer’s visual details matches with the image, we also include the CLIPScore metric Hessel et al. (2021). Including the CLIPScore metric has the added benefit of including an assessment of how close a model-generated answer is to a visual description.
Experimental Setup
For all models, we ran baseline experiments where only the image and question were provided, and the full contextual condition where the scenario and description were additionally provided, just like for the participants who answered the questions. We include our prompt for the contextual condition in Appendix A.1. All evaluations were conducted with the greedy decoding method to ensure reproducibility.
Overall Results
In Table 1, we show the performance of all models when provided with the image and question (baseline) and the full context that was available to human answerers (contextual).
The IDEFICS model with the contextual condition has the highest performance for all models across all metrics except CLIPScore, which remains unaffected. Evidently, integrating contextual information made the content itself more aligned with the ground-truth answer references, as evaluated by the reference-based metrics. The unchanged CLIPScore rating suggests that the contextually induced changes in the generated answers don’t change how they relate to the images.
While for IDEFICS (the best-performing model), CLIPScore is the only metric that doesn’t significantly improve in the contextual condition, this pattern is reversed for all other models. Here, the contextual integration primarily boosts CLIPScore (i.e., the similarity to the image) but results in worse performance according to the reference-based metrics. One potential explanation for this pattern is that the less well-performing models closely reiterate the visual information they receive from the description in the contextual condition, instead of adding new information to answer the question. This would result in higher CLIPScores but lower similarity with the ground-truth answers.
To investigate how much visual information models re-iterate within their generated answers, we compared the average cosine similarity (using Sentence-BERT Reimers and Gurevych (2019) embeddings) between the description and the model’s generated answers. Figure 4 shows that the similarity between the description and the generated answers significantly increases with the contextual condition across models, according to a two-sample t-test analysis. However, the increase in description similarity between the baseline and contextual condition is the lowest for IDEFICS. We conclude that in their answers, the contextual versions of LLaVA, mPLUG-Owl, and BLIP-2 mimic the description more than IDEFICS does, which explains the CLIPScore increase across the baseline and contextual conditions for these three models. These findings suggest that those models might overly emphasize the descriptive details when available, leading to less alignment with the ground-truth answers.
5.1 Quantifying Hallucinations
While IDEFICS (contextual) best approximates the human answers, it’s still far from achieving human-like performance. When a model is intended to make information available that can’t be verified by a user, it is especially important that this model only generates truthful content MacLeod et al. (2017). While this is straightforward to evaluate for accuracy-based VQA datasets, this isn’t easily captured by the similarity-based metrics that are used for long-form evaluation, as presented here. In this section, we aim to give a brief intuition about the rate of hallucinated or incorrect pieces of information contained in the answers of the best-performing model, IDEFICS (contextual).
To estimate how many answers contain wrong, hallucinated or unverifiable content generated by the model, three authors independently annotated a randomly sampled subset of 51 generated answers, with a Fleiss inter-annotator kappa agreement of 49.4%. The annotations show that 38.2% of the model-generated answers contained clearly erroneous information and for another 15%, the truthfulness couldn’t be established. These results indicate that even the best-performing model (IDEFICS (contextual)) is still far from human-like performance on CommVQA.
Since named entities are a likely source of error and hallucination for VLMs, we further analyze the potential accuracy of named entity recognition within the dataset at scale. Without the context, only 35.4% of the unique named entities in IDEFICS’ answer also appeared in any of the reference answers, suggesting that IDEFICS is frequently naming incorrect entities. Providing IDEFICS with the context guides the model to identify the named entities in the answers more accurately, at a rate of 51%. Most of the named entities that IDEFICS (contextual) struggles on are numbers, particularly in counting. Of the misidentified named entities, 38% of the misidentified named entities were cardinals, ordinals, and quantity estimations, and 25% were time and date indicators, indicating that this model still needs to improve in counting and quantity estimations.
Taken together, these results highlight that CommVQA poses a challenging problem for state-of-the-art models. Our analyses suggest that models might not be able to leverage contextual information effectively and the high degree of hallucinations makes them unreliable, highlighting two important areas for future research.
Figure 4: Average cosine similarity to each description from the human-elicited answers and model-generated answers across all conditions. Asterisks indicate significance levels according to a two-sample t-test analysis.
6 Conclusion
Visual question answering models are a promising tool for making visual content accessible to all. With CommVQA, we move the contextually isolated VQA task into a communicative setting that starts reflecting the diversity of downstream use, while keeping close control over the nature of the dataset to ease interpretability. We find strong evidence from dataset analysis that the types of relevant questions and answers change with the contextual domains where images appear. We also find evidence that integrating contextually relevant information improves model performance. Our results suggest that the path towards building viable VQA systems requires a focus on the wider communicative context where images appear.
Limitations
In this work, we show that the scenarios images are presented in fundamentally affect the VQA task. To investigate this, we varied the broad type of website where we embedded the image (Social Media, Shopping, etc). However, context effects are likely much more diverse than the effects studied in this work. For example, recent work suggests that even topic changes within a website domain (a Wikipedia article on Mountains vs. Body of Water) change the information needs that sighted and BLV users have for image descriptions Kreiss et al. (2022b). This result likely translates to the VQA task and needs further investigation.
To allow for easy manipulations of the context domains and a highly controlled recruitment, participants were put in simulated scenarios where they were told about the website domain where the image appears. While even in these induced contextual setups, we find significant contextual variations, future work needs to explore the way this extends to real-world user experience outside of simulated scenarios.
In CommVQA, we elicited questions and answers from sighted participants. Evidence from prior work Stangl et al. (2021) indicates that these results would likely transfer more directly to the accessibility scenario, but future work needs to analyze how the sighted user behavior translates to the BLV population more directly.
Acknowledgements
This research is supported in part by grants from Google and the Generative AI for the Future of Learning program at Stanford. We thank our experiment participants for their invaluable input. We also thank Bay Foley-Cox for his thorough feedback on previous drafts, and all members of the Coalas Lab.
Appendix A Appendix
A.1 Dataset Collection Overview
In this section, we provide a full list of the tasks and prompts provided in each step of the dataset collection process.
For the context matching stage, we prompted GPT-4V with this prompt:
Imagine you are a person browsing the Internet. Please rank the following scenarios in which you might encounter this image:
1) You are browsing a shopping website, with the goal of purchasing an item or experience.
2) You are browsing science magazines (such as National Geographic), with the goal of learning more about recent science developments.
3) You are browsing news websites (such as New York Times), with the goal of learning more about recent news developments.
4) You are browsing a health website, with the goal of learning how to live a healthier lifestyle.
5) You are browsing social media, with the goal of learning more about your connections.
6) You are browsing a travel website, with the goal of traveling to a new location.
For the description elicitation, we asked participants to: “Please edit this description of the image to correct any errors and make the description more useful to someone who cannot see it.” This phrase was intended to incentivize people to both fix errors and include any communicative details that they felt was missing.
For the question elicitation study, we asked participants: “Imagine you are browsing a {scenario} website when you encounter the following image: {description}. If you encounter this image on a {scenario} website, what are two questions you’d want to have answered by someone who can see the image?”
For the answer elicitation study, we asked participants to: “Imagine you are browsing a {scenario} website when you encounter the following image. {image} Another user cannot see the image directly but has access to the following image description: {description}. Based on the description, they asked a follow-up question. Please answer based on the image. {question}.”
We integrated context by prompting models with the format:
Assume someone is browsing a {scenario} website when they encounter this image. They cannot see the image directly, but they can access this image description: {description}. Based on this description, they asked this follow-up question. Please answer based on the image. In your answer, prioritize details relevant to this person. Question: {question}
A.2 Additional Dataset Composition Analysis
Figure 5: Overall named entity recognition analysis on answers written by participants. Error bars represent 95% confidence intervals.
A.2.1 Answer analysis of named entities
As reported for the analysis of the questions (Section 4.2), we find that there is also significant variation in the number of named entities referenced in the answers. As Figure 5 illustrates, most organizations and people are referenced in the news context, and the least are mentioned in the health and social media scenarios, highlighting how the contextual variance in the questions trickles down to the answer distributions.
Figure 6: Example entries in the CommVQA dataset.
A.2.2 Answer analysis of polar questions
We find that participants readily volunteer information that are not strictly necessary for answering the question directly. For this purpose, polar questions in VQA are an interesting case study, since for polar questions, only an answer of “Yes” or “No” is strictly required. However, in CommVQA, the average length of answers to polar questions was 52.1 characters. A direct string matching analysis revealed that only 0.01% of the answers to polar questions were simply “Yes” or “No,” indicating that most of the time participants provided further information. For example, the question “Is this something you think I could do with limited sight?” received the answer “You could possibly do this with limited sight, but I wouldn’t suggest it.” Prior work suggests that such strictly overinformative behavior is driven by considerations of what information is relevant to a questioner’s goals Tsvilodub et al. (2023). Consequently, the fact that we’re seeing this behavior could again be attributed to the communicative setup of the task, and suggests that positioning VQA as such results in answer patterns that explicitly consider relevancy to a questioner’s goals.
However, an answer longer than “Yes” or “No” might be only restating what was in the question–e.g., if the question was “Are there any trees in the image?”, an answer might be, “Yes, there are trees in the image.” As an exploratory analysis, we prompted GPT-4 OpenAI (2023a) to classify whether the provided answer strictly adhered to the scope of the question, or whether it went beyond and offered additional details. Surprisingly, we found that out of the 528 polar questions in our dataset, only 34% stopped at providing enough information to answer the question, while the other 63% answers were classified as providing information beyond what was necessary to answer the question.
Of the answers to polar questions that over-informed, after we conducted further GPT-4 OpenAI (2023a) analysis, approximately 39% were describing visual details not directly asked about in the question (e.g., Question: Are there any other objects in the image? Answer: Small fish) while 60% were expressing uncertainty and making an educated guess or inference (e.g., Question: Is the athlete a professional? Answer: It’s not explicitly clear whether or not the athlete is a professional although one can infer that the athlete is an enthusiast who clearly trains for the sport and enjoys participating in competitions).
A.3 Models Chosen and Parameters
For running model evaluation, we used the following model versions:
  1. 1.
    LlaVA: llava-v1.5-13b-3GB
  2. 2.
    IDEFICS: idefics-9b-instruct
  3. 3.
    mPLUG-Owl: mplug-owl-llama-7b.
  4. 4.
    BLIP-2: blip2-opt-2.7b.